a rocket is fired from the ground with the inital velocity of 110 ft per second. The height h(x) of the rocket at any time, x (in seconds) is given by the function h(x)= -16x^2+110x. Approximatley what is the maxium height of the rocket
Answer by lwsshak3(11628)

a rocket is fired from the ground with the inital velocity of 110 ft per second. The height h(x) of the rocket at any time, x (in seconds) is given by the function h(x)= -16x^2+110x. Approximatley what is the maxium height of the rocket
**
h(x)= -16x^2+110x
This is an equation of a parabola that opens downwards.
complete the square:
h(x)= -16(x^2-110x/16+(55/16)^2)+(55/16)^2 *16
h(x)= -16(x-(55/16)^2+(55/16)^2 *16
h(x)= -16(x-3.44)^2+189.1
maxium height of the rocket≈189.1 ft